OpenStudy (anonymous):

Eric has a gift card to his favorite restaurant. Every Tuesday he has the same meal at the restaurant. What is the equation that represents Eric’s gift card balance? (y – 100) = (x – ) Eric’s Gift Card Weeks 2 4 6 8 Balance $125 $100 $75 $50

OpenStudy (anonymous):

@DanJS @ganeshie8

OpenStudy (danjs):

use 2 of the points to find the slope, then use any of the 4 points as your (x,y) value

OpenStudy (danjs):

i take that back, they gave you y= 100 so the x value is 4

OpenStudy (anonymous):

so the anwer is 4

OpenStudy (danjs):

slope m = (100-25)/(4-2) = 75/2 point(x,y) = (4,100) (y-100) = (75/2)(x-4)

OpenStudy (anonymous):

(y-100) = (75/2)(x-4) thats the anwer??

OpenStudy (danjs):

sorry,

OpenStudy (danjs):

m = (100-125)/(4-2) = -25/2

OpenStudy (anonymous):

thanks